40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5745     Entry time: Thu Oct 27 03:32:45 2011
Author: Koji 
Type: Summary 
Category: IOO 
Subject: RFAM monitor progress 

[Suresh, Mirko, Koji]

A cable from the stochmon box to the cross connect for the EPICS ADCs is installed.

The power supply and the signal outputs are concentrated in a single DSub 9pin connector
that is newly attached on the box.

The connection from the stochmon side of the cable and the EPICS value was confirmed.
The calibration of them looks fine.

To do:
- Once the stochmon box is completed we can immediately test it.
- The EPICS channel names are still as they were. We need to update the database file of c1iool0, the chans file for the slow channel.


The pinout is as following

-------------
| 1 2 3 4 5 |   Female / Inside View
\  6 7 8 9  /
 \---------/

1 - 11MHz Signal
2 - 30MHz Signal
3 - 55MHz Signal
4 - NC
5 - +5V supply
6 - 11MHz Return
7 - 30MHz Return
8 - 55MHz Return
9 - Supply ground

ELOG V3.1.3-